OpenStudy (anonymous):

Implicit Differentiation: x + cos (x+y) = 0 A) csc (x+y)-1 B) csc (x+y) C) 1/(sin(x+y)) D) 1/(1-x^2)^.5 E) 1-sin(x)/sin(y) I did the question twice, not getting any of the answers listed.

OpenStudy (kainui):

Show me what you did and I'll show you where you went wrong and help you out. =D

OpenStudy (kainui):

I have the right answer here, I know how to do it, but I can't help you if you don't work with me.

OpenStudy (anonymous):

\[x+\cos(x+y)=0\] So I took the derivative (implicitly), and my first step ended up with... \[1-\sin(x+y)*1+y'\] I think my confusion is stemming from the inside part of the sine function, I know I chain rule, but did I differentiate the inside correctly?

OpenStudy (kainui):

You did the differentiation correctly, but remember that it is (1+y') in parentheses and it's all equal to 0. Now you just need to solve for y'.

OpenStudy (anonymous):

So by simplification... \[1-\sin(x+y)*(1+y')=0\] \[\sin(x+y)*(1+y')=-1\] \[(1+y')=\frac{-1}{sin(x+y)}\] \[y'=\frac{-1}{sin(x+y)}-1\] Or did I do something wrong on the way there?

OpenStudy (kainui):

You should have 1=sin(x+y)(1+y'), not a negative one. Also, remember that 1/sin(x)=csc(x)